the value of specific heat for copper is

K) 12 of 21 What will be the equibrium emperature when a 275 g block of copper at 255 Cis placed in a 145 g Constants Express your answer using three significant figures Tho value of 390 J/kg. C, for aluminun is 000 J/kg.O,and for water is 4186 J/kg C heat for copper is Previous Answers
Answer #1

Let us assume that the final equilibrium temperature is T. Then, we have from the colorimetry,
  Leat lost by the copper = heat gain by the ( water + aluminium )
   \Rightarrow m_cc_c(t_c-T)=m_ac_a(T-t_a)+m_wc_w(T-t_w)
   \Rightarrow T=\frac{m_cc_ct_c+m_ac_at_a+m_wc_wt_w}{m_cc_c+m_ac_a+m_wc_w}
And putting the given values,
   mc 275 g, ma -145 g, my 815 q
   c_c=390~J/kg.^\circ C,~~c_a=900~J/kg.^\circ C,~~c_w=4186~J/kg.^\circ C
And
t_c=255~^\circ C,~~t_a=t_w=14.0^\circ C
And so, we have
\Rightarrow T=\frac{(275\times 390\times 255)+(145\times 900\times 14)+(815\times 4186\times 14)}{(275\times 390)+(145\times 900)+(815\times 4186)}
  \Rightarrow T=21.08~\degree C